# Causal Classification of Pathological Misner-Type Spacetimes

### v4

@@ Abstract.
We investigate three causality-violating spacetimes: Misner space (including Kip Thorne’s “moving wall” model), the pseudo-Schwarzschild spacetime, and a new model introduced here, the pseudo-Reissner-Nordström spacetime. Despite their different physical origins—ranging from a flat vacuum solution to a black-hole-type vacuum solution to a non-vacuum model requiring exotic matter—all three share a common warped-product structure, 22–dimensional cylindrical base metrics of Eddington-Finkelstein type, and fundamental causal features such as Cauchy and chronology horizons, acausal regions, and analogous geodesic behaviour. Building on a conjecture first proposed in 2016, we present a formal proof that the three models are pairwise isocausal on their universal covers and on suitable causally regular regions of their compactified forms. The proof is constructive, providing explicit causal bijections on the covers and identifying a concrete deck-equivariance criterion governing descent to the compactified spacetimes: if the equivariance degree satisfies $|k|=1$ the models are globally isocausal, whereas if $|k|>1$ or equivariance fails, then at most a one-way causal relation holds between the compactified models. These results supply a rigorous causal classification linking these spacetimes, placing them within a unified Misner-type family and providing a framework for extending the classification to other causality-violating solutions.
@@
of a project during 2016-2018 under the supervision of Kip S. Thorne.
@@ 1. Introduction
General relativity defines a class of cosmological models, each representing an idealization of a physically possible universe compatible with the theory. A cosmological model is a mathematical description of this idealized universe, referred to as a spacetime and represented by a pair $(M,g)$, where $M$ is a differentiable manifold capturing the topology and continuity of the universe, and $g$ is a smooth, symmetric, non-degenerate (0,2)(0,2)–tensor field encoding the geometric and causal structure. Each point in $M$ represents an event, and the causal structure determined by $g$ governs the possible trajectories of particles and light rays. Within this framework, the theoretical possibility of time travel—through closed timelike curves (CTCs) that allow a timelike observer to return to an event in their own past—has been studied extensively [1, 4, 9, 19, 12, 21, 22]. In 1967, Charles Misner introduced the Misner space [14] as a minimal example of a spacetime with a chronology-violating region. While constructing spacetimes containing CTCs is relatively straightforward, finding examples with a plausible physical interpretation is more challenging. The pseudo-Schwarzschild spacetime, proposed by Ori [16] in 2007, is a more sophisticated causality-violating model that incorporates features such as Cauchy horizons, avoids singularities, and models the breakdown of determinism inside black holes. [^1] An “idealized black hole model” is a simplified theoretical depiction of a black hole, commonly assuming perfect spherical symmetry and no rotation, and typically derived from the Schwarzschild metric.
@@
We then introduce a new chronology-violating spacetime: the pseudo-Reissner-Nordström spacetime, which generalizes the previous two by including an electric charge parameter in analogy with the classical Reissner-Nordström solution. Like its predecessors, this model possesses Cauchy and chronology horizons, acausal regions, and a warped-product structure with a 22–dimensional cylindrical base and hyperbolic spatial fibers. Its inclusion is especially noteworthy because it is a non-vacuum solution, in contrast to the vacuum nature of Misner and pseudo-Schwarzschild spacetimes, yet it still exhibits the same characteristic causal pathologies.
@@
From a mathematical perspective, the shared warped-product structure and conformal relation of the base metrics make these isocausality results natural; from a physical perspective, the fact that such structurally and interpretatively different spacetimes exhibit identical causal behaviour is striking. This work provides a unified geometric and causal framework for understanding these examples, and paves the way for extending the classification to other models—such as the pseudo-Kerr spacetime—that may also belong to this Misner-type, causality-violating family.
@@ 2.1. Kip Thorne’s physics approach to Misner space
Consider the flat 22–dimensional Minkowski spacetime with a reference frame in Lorentzian coordinates $(t,x)$, where $t$ is interpreted as the time direction and $x$ as the space direction. Next, we identify the tt–axis at $x=0$ with the tt–axis at $x=P$ (which we will henceforth refer to as the $\bar{t}$–axis): $(t,P)\sim(t,0)$. At $t=0$, we set the $\bar{t}$–axis into motion with constant speed $\beta$ towards the tt–axis, while the reference frame in Lorentzian coordinates remains at rest with respect to the tt–axis. We can think of the tt–axis as representing an observer at rest and the $\bar{t}$–axis as representing an observer moving at speed $\beta$ relative the reference frame. This moving observer is equipped with a clock, and $\tau$ denotes the proper time measured by this clock. [^6] Proper time, also called clock time or process time, is a measure of the amount of physical process that a system undergoes. This x,tx,t–coordinate system is furnished with the flat Minkowski metric $\widetilde{\eta}$, and therefore the light cones do not tilt or open. Due to the special relativistic time dilation, a region with closed timelike curves develops, see Figure 1.
Note that the worldline of the clock at rest (along the tt–axis) and the worldline of the clock in motion (along the $\bar{t}$–axis) are identified by $\sim$, meaning the events along these lines are physically identical. However, due to special relativity, the observer at rest perceives the moving clock to tick more slowly. On the tt–axis we have $\tau=t$, while on the $\bar{t}$–axis, due to time dilation, $\tau=\frac{t}{\sqrt{1-\beta^{2}}}=\gamma t$, where $\gamma:=\frac{1}{\sqrt{1-\frac{v^{2}}{c^{2}}}}=\frac{1}{\sqrt{1-\beta^{2}}}$.
@@
Figure 2. ‘Moving wall’ Misner space with proper time origins, such that the two walls are separated by the chronology horizon at $x=t$ by the spatial distance $D$.
@@
where $\xi=\sqrt{\left(\frac{1-\beta}{1+\beta}\right)}$ is the Doppler blueshift. In the covering space (Figure 3) we get the spatial distance between $W_{0}$ and the $\tilde{t}$–axis by the converging series
@@
To summarize, this ensures that the intersection points $s_{n}$ are not physical singularities, but merely coordinate singularities. This reasoning is supported by the fact that the 22–dimensional Misner universe can be viewed as a simplified $2D$ wormhole spacetime [12]; both spacetimes share the same covering space. By projecting a wormhole whose mouths move past each other at speed $\beta$ [21, Figure 3c] onto the xtx--t plane, we obtain a situation similar to that of Misner space, as described above. However, this wormhole spacetime diagram reveals a new feature: a past chronology horizon, $\mathcal{H}_{-}$. Therefore, the chronology-violating region is confined by two chronology horizons. Like the future chronology horizon $\mathcal{H}_{+}$, the past chronology horizon $\mathcal{H}_{-}$ is a null surface.
@@ 2.2. Geodesics for warped Misner product space
The 44–dimensional Misner space (2.3) can be expressed as a warped product space of the form $M=M_{Z}\times_{r}H^{2}$, where $M_{Z}$ is the 22–dimensional Misner-type spacetime with coordinates $(T,\varphi)$, and $H^{2}$ is the 22–dimensional hyperbolic plane, described by coordinates $(\theta,\phi)$. The general form of the metric combines the 22–dimensional Misner-type geometry, which encodes the non-trivial topological and causal structure, with a 22–hyperbolic plane that introduces hyperbolic symmetry. The metric can be written as
@@
where $f(T)$ describes the time-dependent geometry of the Misner-type spacetime and $r$ is the warping function. To interpret this, we consider $\mathbb{R}^{2}$ as being represented topologically by the hyperbolic plane $H^{2}$. The 44–dimensional Misner space can then be described as a warped product $M:=M_{Z}\times_{r}H^{2}$, with the warping function $r$ being a scalar function $r:M_{Z}\rightarrow\mathbb{R}$, depending on the coordinates of the base manifold $M_{Z}$. The Misner space $M_{Z}=S^{1}\times\mathbb{R}$ is equipped with the cylindrical metric $ds^{2}=2dTd\varphi+Td\varphi{}^{2}$. The fiber $H^{2}$, with coordinates $(\theta,\phi)$, has the hyperbolic metric $g_{H^{2}}=d\theta^{2}+\sinh^{2}\theta d\phi^{2}$, scaled by $r^{2}$. Therefore, the metric on the total space $M$ can be written as $ds^{2}=g_{Z}+r^{2}g_{H^{2}}$, which explicitly expands to
@@
Based on the warped geometry of Misner space, we can focus on the embedded cylinders defined by $\theta=\textrm{const}$ and $\varphi=\textrm{const}$. By doing so, we restrict our attention to the 22–dimensional Misner cylinder $M_{Z}$, which captures the essential features of the Misner-type geometry.
@@
Given the 22–dimensional Misner metric $ds^{2}=Td\varphi{}^{2}+2dTd\varphi$, null geodesics are described by
@@ 3. General equations for radial geodesics
In this section, we present a complete canonical formulation of the geodesic equations for cylindrical Misner-type spacetime models, focusing on the embedded $\theta=\textrm{const}$ and $\varphi=\textrm{const}$ cylinders, as introduced in Subsection 2.2. We anticipate that the radial geodesic equations for the pseudo-Schwarzschild and pseudo-Reissner-Nordström spacetimes (to be discussed in Sections 4 and 6) will exhibit structural similarities. To unify these cases, we derive a general form of the geodesic equations, which proves to be valid for all three hyperbolically symmetric spacetimes. Notably, this general form can also be applied to Misner space, recognizing that the 22–dimensional Misner metric, $ds^{2}=Td\varphi{}^{2}+2dTd\varphi$, can be rewritten in the canonical form
@@
To generalize the Equation 3.1 to the 44–dimensional case for hyperbolically symmetric models, we start with the general form of a hyperbolically symmetric metric given by
@@
These results provide further insights. By renaming the coordinates as $\nu=\psi$ and $r=T$, we can carry over the findings to the Misner case, which takes the same general form as the 22–dimensional cylindrical pseudo-Schwarzschild (4.4) and pseudo-Reissner-Nordström (6.3) spacetimes in Eddington-Finkelstein coordinates:
@@ 4. Pseudo-Schwarzschild spacetime
Although the pseudo-Schwarzschild spacetime is derived from the well-known Schwarzschild spacetime, its global and causal structure more closely resembles that of Misner space. The pseudo-Schwarzschild solution satisfies the Einstein field equations in vacuum and describes a static, hyperbolically symmetric vacuum spacetime [16]. Furthermore, its universal covering space asymptotically approaches Minkowski space, which is known as asymptotic flatness. This serves as another example of a spacetime that violates the chronology condition and allows for the existence of closed timelike curves. The geometry of the pseudo-Schwarzschild spacetime admits a warped product structure, with the base manifold $S^{1}\times\mathbb{R}^{+}$ and the hyperbolic plane $H^{2}$ as the fiber. The point set that defines the pseudo-Schwarzschild manifold is
@@
where $S^{1}$ is the 11–sphere, topologically equivalent to an interval whose endpoints are glued together by identification, and $H_{r}^{2}$ (when $t$, $r=\textrm{const}$) represents a 22–dimensional surface that can be identified with one of the sheets of the two-sheeted space-like hyperboloid. [^7] The upper sheet of the hyperboloid can be globally embedded in $3$-dimensional Minkowski space. We can express the upper sheet hyperboloid using a parameterization in polar coordinates $(\theta,\phi)$. Specifically, the map $\mathbb{R}^{+}\times S^{1}\rightarrow H_{r}^{2}$, $(\theta,\phi)\mapsto(r{\normalcolor\cosh(\theta)},r{\normalcolor\sinh(\theta)}{\normalcolor\cos(\varphi)},r{\normalcolor\sinh(\theta)\sin(\phi)})$ describes the embedding of the upper sheet of the hyperboloid in Minkowski space. The hyperboloid together with the induced Riemannian metric $g=d\theta\otimes d\theta+\sinh^{2}(\theta)d\phi\otimes d\phi$ is called hyperbolic plane. Then the pseudo-Schwarzschild metric can be derived by performing a Wick rotation $\theta\rightarrow i\theta$ and a signature change from the famous Schwarzschild metric as follows:
@@
The pseudo-Schwarzschild spacetime is static inside the horizon and time-dependent outside of the horizon. For $r<2m$, the tt–direction $\partial_{t}$ is timelike, and $r$ is a spatial coordinate and the $r$-direction $\partial_{r}$ is spacelike, as seen from the metric components:
@@
Recalling that the pseudo-Schwarzschild spacetime is $M=M_{Z}\times_{r}H^{2}$, and its metric can be decomposed into the cylindrical metric $g_{Z}$ and the hyperbolic metric $g_{H^{2}}$, we proceed by setting $\theta=\textrm{const}$ and $\phi=\textrm{const}$. Allowing $\nu$ and $r$ to vary and take the values as previously defined, we obtain the 22–dimensional cylindrical pseudo-Schwarzschild spacetime $M_{Z}=S^{1}\times\mathbb{R}^{+}$. This spacetime is described by the degenerate metric in Eddington-Finkelstein coordinates (see Figure 6):
(4.4) $g_{Z}:=-(\frac{2m}{r}-1)d\nu^{2}+2d\nu dr.$
In two dimensions, the Christoffel symbols are identical for both the Schwarzschild and pseudo-Schwarzschild spacetimes. Therefore, the Riemann curvature tensor for the cylindrical pseudo-Schwarzschild spacetime (4.3) is given by $R_{trtr}=-\frac{2m}{r^{3}}\eqqcolon K=\frac{1}{2}S$, where $K$ denotes the Gaussian curvature and $S$ represents the scalar curvature. Since we require $m$ to be non-negative and $r$ to be positive, the Gaussian curvature is non-positive throughout the cylinder. [^9] The Riemannian curvature tensor for the 44–dimensional pseudo-Schwarzschild spacetime can easily computed by the prescription of Wick rotation and sign change.

### v5

@@ Abstract.
We investigate three causality-violating spacetimes: Misner space (including Kip Thorne’s “moving wall” model), the pseudo-Schwarzschild spacetime, and a new model introduced here, the pseudo-Reissner-Nordström spacetime. Despite their different physical origins—ranging from a flat vacuum solution to a black-hole-type vacuum solution to a non-vacuum model requiring exotic matter—all three share a common warped-product structure, $2$-dimensional cylindrical base metrics of Eddington-Finkelstein type, and fundamental causal features such as Cauchy and chronology horizons, acausal regions, and analogous geodesic behaviour. Building on a conjecture first proposed in 2016, we present a formal proof that the three models are pairwise isocausal on their universal covers and on suitable causally regular regions of their compactified forms. The proof is constructive, providing explicit causal bijections on the covers and identifying a concrete deck-equivariance criterion governing descent to the compactified spacetimes: if the equivariance degree satisfies $|k|=1$ the models are globally isocausal, whereas if $|k|>1$ or equivariance fails, then at most a one-way causal relation holds between the compactified models. These results supply a rigorous causal classification linking these spacetimes, placing them within a unified Misner-type family and providing a framework for extending the classification to other causality-violating solutions.
@@
of a project during 2016–2018 under the supervision of Kip S. Thorne.
@@ 1. Introduction
General relativity defines a class of cosmological models, each representing an idealization of a physically possible universe compatible with the theory. A cosmological model is a mathematical description of this idealized universe, referred to as a spacetime and represented by a pair $(M,g)$, where $M$ is a differentiable manifold capturing the topology and continuity of the universe, and $g$ is a smooth, symmetric, non-degenerate $(0,2)$-tensor field encoding the geometric and causal structure. Each point in $M$ represents an event, and the causal structure determined by $g$ governs the possible trajectories of particles and light rays. Within this framework, the theoretical possibility of time travel—through closed timelike curves (CTCs) that allow a timelike observer to return to an event in their own past—has been studied extensively [1, 4, 9, 12, 19, 21, 22]. In 1967, Charles Misner introduced the Misner space [14] as a minimal example of a spacetime with a chronology-violating region. While constructing spacetimes containing CTCs is relatively straightforward, finding examples with a plausible physical interpretation is more challenging. The pseudo-Schwarzschild spacetime, proposed by Ori [16] in 2007, is a more sophisticated causality-violating model that incorporates features such as Cauchy horizons, avoids singularities, and models the breakdown of determinism inside black holes. [^1] An “idealized black hole model” is a simplified theoretical depiction of a black hole, commonly assuming perfect spherical symmetry and no rotation, and typically derived from the Schwarzschild metric.
@@
We then introduce a new chronology-violating spacetime: the pseudo-Reissner-Nordström spacetime, which generalizes the previous two by including an electric charge parameter in analogy with the classical Reissner-Nordström solution. Like its predecessors, this model possesses Cauchy and chronology horizons, acausal regions, and a warped-product structure with a $2$-dimensional cylindrical base and hyperbolic spatial fibers. Its inclusion is especially noteworthy because it is a non-vacuum solution, in contrast to the vacuum nature of Misner and pseudo-Schwarzschild spacetimes, yet it still exhibits the same characteristic causal pathologies.
While the spacetimes considered in this work are mathematically well-defined solutions of Einstein’s equations, some of them—most notably the pseudo-Reissner-Nordström model—require violations of classical energy conditions. Consequently, their direct physical realizability in a classical astrophysical context is a subject of ongoing debate. Nevertheless, the investigation of such causality-violating geometries is well motivated for several reasons. First, they serve as controlled testbeds for probing the limits of general relativity, particularly in regimes where classical notions of causality and determinism break down. Second, these models provide insight into the structure and stability of Cauchy horizons, which are central features of physically relevant solutions such as the interior of charged or rotating black holes. Third, the study of spacetimes admitting closed timelike curves is closely related to foundational questions, including the validity of the chronology protection conjecture proposed by Stephen Hawking, and the role of quantum effects in enforcing causal consistency. In particular, such examples help clarify the role of classical energy conditions by exhibiting the types of pathological behavior—such as closed timelike curves or breakdowns of predictability—that these conditions are intended to exclude. From this perspective, even spacetimes supported by exotic matter can illuminate which features of causality violation are generic and which depend sensitively on the underlying matter content. The unified framework developed in this work therefore contributes not only to the classification of specific models, but also to a broader understanding of the interplay between geometry, causality, and physical viability in general relativity.
@@
From a mathematical perspective, the shared warped-product structure and conformal relation of the base metrics make these isocausality results natural; from a physical perspective, they suggest that key causal features may arise independently of the detailed matter content. This work provides a unified geometric and causal framework for understanding these examples, and paves the way for extending the classification to other models—such as the pseudo-Kerr spacetime—that may also belong to this Misner-type, causality-violating family.
@@ 2.1. Kip Thorne’s physics approach to Misner space
Consider the flat $2$-dimensional Minkowski spacetime with a reference frame in Lorentzian coordinates $(t,x)$, where $t$ is interpreted as the time direction and $x$ as the space direction. Next, we identify the $t$-axis at $x=0$ with the $t$-axis at $x=P$ (which we will henceforth refer to as the $\bar{t}$-axis): $(t,P)\sim(t,0)$. At $t=0$, we set the $\bar{t}$-axis into motion with constant speed $\beta$ towards the $t$-axis, while the reference frame in Lorentzian coordinates remains at rest with respect to the $t$-axis. We can think of the $t$-axis as representing an observer at rest and the $\bar{t}$-axis as representing an observer moving at speed $\beta$ relative the reference frame. This moving observer is equipped with a clock, and $\tau$ denotes the proper time measured by this clock. [^6] Proper time, also called clock time or process time, is a measure of the amount of physical process that a system undergoes. This $x,t$-coordinate system is furnished with the flat Minkowski metric $\widetilde{\eta}$, and therefore the light cones do not tilt or open. Due to the special relativistic time dilation, a region with closed timelike curves develops, see Figure 1.
@@
Note that the worldline of the clock at rest (along the $t$-axis) and the worldline of the clock in motion (along the $\bar{t}$-axis) are identified by $\sim$, meaning the events along these lines are physically identical. However, due to special relativity, the observer at rest perceives the moving clock to tick more slowly. On the $t$-axis we have $\tau=t$, while on the $\bar{t}$-axis, due to time dilation, $\tau=\frac{t}{\sqrt{1-\beta^{2}}}=\gamma t$, where $\gamma:=\frac{1}{\sqrt{1-\frac{v^{2}}{c^{2}}}}=\frac{1}{\sqrt{1-\beta^{2}}}$.
@@
Figure 2. “Moving wall” Misner space with proper time origins, such that the two walls are separated by the chronology horizon at $x=t$ by the spatial distance $D$.
@@
where $\xi=\sqrt{\left(\frac{1-\beta}{1+\beta}\right)}$ is the Doppler blueshift.
In the covering space (Figure 3) we get the spatial distance between $W_{0}$ and the $\tilde{t}-$axis by the converging series
@@
To summarize, this ensures that the intersection points $s_{n}$ are not physical singularities, but merely coordinate singularities. This reasoning is supported by the fact that the $2$-dimensional Misner universe can be viewed as a simplified $2D$ wormhole spacetime [12]; both spacetimes share the same covering space. By projecting a wormhole whose mouths move past each other at speed $\beta$ [21, Figure 3c] onto the xtx-t plane, we obtain a situation similar to that of Misner space, as described above. However, this wormhole spacetime diagram reveals a new feature: a past chronology horizon, $\mathcal{H}_{-}$. Therefore, the chronology-violating region is confined by two chronology horizons. Like the future chronology horizon $\mathcal{H}_{+}$, the past chronology horizon $\mathcal{H}_{-}$ is a null surface.
@@ 2.2. Geodesics for warped Misner product space
The $4$-dimensional Misner space (2.3) can be expressed as a warped product space of the form $M=M_{Z}\times_{r}H^{2}$, where $M_{Z}$ is the $2$-dimensional Misner-type spacetime with coordinates $(T,\varphi)$, and $H^{2}$ is the $2$-dimensional hyperbolic plane, described by coordinates $(\theta,\phi)$. The general form of the metric combines the $2$-dimensional Misner-type geometry, which encodes the non-trivial topological and causal structure, with a $2$-hyperbolic plane that introduces hyperbolic symmetry. The metric can be written as
@@
where $f(T)$ describes the time-dependent geometry of the Misner-type spacetime and $r$ is the warping function. To interpret this, we consider $\mathbb{R}^{2}$ as being represented topologically by the hyperbolic plane $H^{2}$. The $4$-dimensional Misner space can then be described as a warped product $M:=M_{Z}\times_{r}H^{2}$, with the warping function $r$ being a scalar function $r:M_{Z}\rightarrow\mathbb{R}$, depending on the coordinates of the base manifold $M_{Z}$. The Misner space $M_{Z}=S^{1}\times\mathbb{R}$ is equipped with the cylindrical metric $ds^{2}=2dTd\varphi+Td\varphi{}^{2}$. The fiber $H^{2}$, with coordinates $(\theta,\phi)$, has the hyperbolic metric $g_{H^{2}}=d\theta^{2}+\sinh^{2}\theta d\phi^{2}$, scaled by $r^{2}$. Therefore, the metric on the total space $M$ can be written as $ds^{2}=g_{Z}+r^{2}g_{H^{2}}$, which explicitly expands to
@@
Based on the warped geometry of Misner space, we can focus on the embedded cylinders defined by $\theta=\textrm{const}$ and $\varphi=\textrm{const}$. By doing so, we restrict our attention to the $2$-dimensional Misner cylinder $M_{Z}$, which captures the essential features of the Misner-type geometry.
@@
Given the $2$-dimensional Misner metric $ds^{2}=Td\varphi{}^{2}+2dTd\varphi$, null geodesics are described by
@@ 3. General equations for radial geodesics
In this section, we present a complete canonical formulation of the geodesic equations for cylindrical Misner-type spacetime models, focusing on the embedded $\theta=\textrm{const}$ and $\varphi=\textrm{const}$ cylinders, as introduced in Subsection 2.2. We anticipate that the radial geodesic equations for the pseudo-Schwarzschild and pseudo-Reissner-Nordström spacetimes (to be discussed in Sections 4 and 6) will exhibit structural similarities. To unify these cases, we derive a general form of the geodesic equations, which proves to be valid for all three hyperbolically symmetric spacetimes. Notably, this general form can also be applied to Misner space, recognizing that the $2$-dimensional Misner metric, $ds^{2}=Td\varphi{}^{2}+2dTd\varphi$, can be rewritten in the canonical form
@@
To generalize the Equation 3.1 to the $4$-dimensional case for hyperbolically symmetric models, we start with the general form of a hyperbolically symmetric metric given by
@@
These results provide further insights. By renaming the coordinates as $\nu=\psi$ and $r=T$, we can carry over the findings to the Misner case, which takes the same general form as the $2$-dimensional cylindrical pseudo-Schwarzschild (4.4) and pseudo-Reissner-Nordström (6.3) spacetimes in Eddington-Finkelstein coordinates:
@@ 4. Pseudo-Schwarzschild spacetime
Although the pseudo-Schwarzschild spacetime is derived from the well-known Schwarzschild spacetime, its global and causal structure more closely resembles that of Misner space. The pseudo-Schwarzschild solution satisfies the Einstein field equations in vacuum and describes a static, hyperbolically symmetric vacuum spacetime [8, 16]. Furthermore, its universal covering space asymptotically approaches Minkowski space, which is known as asymptotic flatness. This serves as another example of a spacetime that violates the chronology condition and allows for the existence of closed timelike curves. The geometry of the pseudo-Schwarzschild spacetime admits a warped product structure, with the base manifold $S^{1}\times\mathbb{R}^{+}$ and the hyperbolic plane $H^{2}$ as the fiber. The point set that defines the pseudo-Schwarzschild manifold is
@@
where $S^{1}$ is the $1$-sphere, topologically equivalent to an interval whose endpoints are glued together by identification, and $H_{r}^{2}$ (when $t$, $r=\textrm{const}$) represents a $2$-dimensional surface that can be identified with one of the sheets of the two-sheeted space-like hyperboloid. [^7] The upper sheet of the hyperboloid can be globally embedded in $3$-dimensional Minkowski space. We can express the upper sheet hyperboloid using a parameterization in polar coordinates $(\theta,\phi)$. Specifically, the map $\mathbb{R}^{+}\times S^{1}\rightarrow H_{r}^{2}$, $(\theta,\phi)\mapsto(r{\normalcolor\cosh(\theta)},r{\normalcolor\sinh(\theta)}{\normalcolor\cos(\varphi)},r{\normalcolor\sinh(\theta)\sin(\phi)})$ describes the embedding of the upper sheet of the hyperboloid in Minkowski space. The hyperboloid together with the induced Riemannian metric $g=d\theta\otimes d\theta+\sinh^{2}(\theta)d\phi\otimes d\phi$ is called hyperbolic plane. Then the pseudo-Schwarzschild metric can be derived by performing a Wick rotation $\theta\rightarrow i\theta$ and a signature change from the famous Schwarzschild metric as follows:
@@
The pseudo-Schwarzschild spacetime is static inside the horizon and time-dependent outside of the horizon. For $r<2m$, the $t$-direction $\partial_{t}$ is timelike, and $r$ is a spatial coordinate and the $r$-direction $\partial_{r}$ is spacelike, as seen from the metric components:
@@
Recalling that the pseudo-Schwarzschild spacetime is $M=M_{Z}\times_{r}H^{2}$, and its metric can be decomposed into the cylindrical metric $g_{Z}$ and the hyperbolic metric $g_{H^{2}}$, we proceed by setting $\theta=\textrm{const}$ and $\phi=\textrm{const}$. Allowing $\nu$ and $r$ to vary and take the values as previously defined, we obtain the $2$-dimensional cylindrical pseudo-Schwarzschild spacetime $M_{Z}=S^{1}\times\mathbb{R}^{+}$. This spacetime is described by the degenerate metric in Eddington-Finkelstein coordinates (see Figure 6):
(4.4) $g_{Z}:=-(\frac{2m}{r}-1)d\nu^{2}+2d\nu dr.$
@@
In two dimensions, the Christoffel symbols are identical for both the Schwarzschild and pseudo-Schwarzschild spacetimes. Therefore, the Riemann curvature tensor for the cylindrical pseudo-Schwarzschild spacetime (4.3) is given by $R_{trtr}=-\frac{2m}{r^{3}}\eqqcolon K=\frac{1}{2}S$, where $K$ denotes the Gaussian curvature and $S$ represents the scalar curvature. Since we require $m$ to be non-negative and $r$ to be positive, the Gaussian curvature is non-positive throughout the cylinder. [^9] The Riemannian curvature tensor for the $4$-dimensional pseudo-Schwarzschild spacetime can easily computed by the prescription of Wick rotation and sign change.